OpenStudy (anonymous):

Which equation would best help solve the following problem? The height of a triangle is 6 m more than its base. The area of the triangle is 56 m2. Find the length of the base. a. 56=1/2(x)(x+6) b. 56=1/2(x2+6) c. 56=x(x+6) d. 56=x2+6

OpenStudy (anonymous):

area = 1/2 x base x perp. height 56 = 1/2x (base)x (base + 6) so a.